Hi can you help me with the following problem?
Write the answer in standard form of scientific notation correctly rounded to 2 figures.
(5.46 x 10^-8) ( 9.77 x 10^-8)
______________________________
(4.07 x 10^3) ( 8.37 x 10^3)
Thanks.
:)

Answer by checkley71(8403)

5.46*9.77=82.6542
10^-8*10^-8=10^-16
THE NUMERATOR=8.26542*10^-15
4.07*8.37=34.0659
10^3*10^3=10^6
THE DENOMINATOR=3.40659*10^7
8.26542*10^-15/3.40659*10^7
2.4263*10^-22